# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v6

@@ 4.3 Constrained Optimization
We next illustrate the scenarios in which the LL cost $L$ in (1b) is subject to the general constraints in (1c). We discuss two cases with equality and inequality constraints, respectively. Constrained optimization is a thoroughly explored field, and related findings were presented in (Dontchev et al., 2009) and summarized in (Gould et al., 2021). This study will focus on the integration of constrained optimization into our special form of BLO (1) under the framework of IL.
@@
[…]
@@
which means that the gradient of $L$ w.r.t. $\bm{\phi}$ has the same direction as the gradient of constraint $\xi$ w.r.t. $\bm{\phi}$ at the optimal solution ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$. Taking the derivative of (39) w.r.t. $\bm{\psi}$ and noting that $\bm{\phi}^{*}$ and $\lambda$ are functions of $\bm{\psi}$, thus
@@
[…]
@@
While we have derived closed-form solutions for the equality constraint in (34) and the inequality constraint in (43), calculating the matrix inversions $H^{-1}$ and $[L_{\bm{\phi}}H^{-1}L_{\bm{\phi}}^{\top}]^{-1}$ can be computationally expensive. To mitigate this, one may employ the approach discussed in Section 4.1, which involves approximating the matrix-inverse-vector product $H^{-1}y$ by solving the linear system $\min_{v}\frac{1}{2}v^{\top}Hv-v^{\top}y$. This complicated computing process can sometimes be simplified by specific robot settings. We next present an example of robot control illustrating this process.
@@
When the Hessian matrices $\frac{\partial^{2}L}{\partial^{2}\phi}$ and $\frac{\partial^{2}\xi}{\partial^{2}\phi}$ of the LL function $L$ and the constraint function $\xi$ are invertible, we can guarantee the matrix $H$ in (37) is invertible. One sufficient condition for such invertibility is that $L$ and $\xi$ are strongly convex. Motivated by the approximate Karush-Kuhn-Tucker (KKT) condition (Andreani et al., 2010), which is characterized as an optimality condition for nonlinear program, we can use the residual function in (Yao et al., 2024b, Eq. (15)) as an optimality measurement. As a result, the optimality can be established and guaranteed under these conditions.
@@
Modeling and control of nonlinear dynamics precisely is critical for robotic applications such as aerial manipulation (He et al., 2023; Geng and Langelaan, 2020) and off-road navigation (Triest et al., 2023). Although classic methods based on physics have excellent generalization abilities, they rely heavily on problem-specific parameter tuning. Methods like state feedback (Mellinger and Kumar, 2011) and optimal control (Garcia et al., 1989; Ji et al., 2022b) require precise system modeling to predict and control dynamics effectively. However, it is challenging to accurately model unpredictable factors such as wind gusts, boundary layer effects, uneven terrain, or hidden dynamics of chaotic nonlinear effects.
@@
In recent years, researchers have spent efforts to integrate physical modeling into data-driven methods and have gained remarkable success in the field of system control (O’Connell et al., 2022; Hu et al., 2023). For example, Amos et al. (2018) developed the differentiable model predictive control (MPC) to combine the physics-based modeling with data-driven methods, enabling learning dynamic models and control policies in an end-to-end manner (Jin et al., 2020). However, many prior studies depend on expert demonstrations or labeled data for supervised learning (Jin et al., 2022; Song and Scaramuzza, 2022). While supervised learning yields effective results in the training datasets, it suffers from challenging conditions such as unseen environments and external disturbances.
@@
This discrepancy between the predicted states and the IMU network measurement reflects the models’s imperfectness, indicating the learnable parameters ${\bm{\theta}}$ in both the network and the MPC module can be adjusted. Specifically, we adopt an IMU denoising model, AirIMU (Qiu et al., 2024) as the perception network, which is a neural model that can propagate system uncertainty through a time horizon. The differentiable MPC (Diff-MPC) from PyPose (Wang et al., 2023) is selected as the physics-based reasoning module.
@@ 4.4 Second-order Optimization
[…]
@@
The theoretical guarantee is similar to that for the first-order methods in Section 4.2, as long as the second-order LL optimizers such as Gauss-Newton and LM approximate $\mu^{*}$ and $\nu^{*}$ to the target accuracy $\epsilon$. For example, if the LL problem is strongly convex, the second-order optimization can find an $\epsilon$-accurate solution with an accelerated linear convergence rate (Ji and Liang, 2021, Theorem 9 and 10).
@@
Simultaneous localization and mapping (SLAM) is a critical technology in computer vision and robotics (Aulinas et al., 2008; Cadena et al., 2016). It aims to simultaneously track the trajectory of a robot and build a map of the environment. SLAM is essential in diverse robotic applications, such as indoor navigation (Wang et al., 2017a), underwater exploration (Suresh et al., 2020), and space missions (Dor et al., 2021). A SLAM system generally adheres to a front-end and back-end architecture. Specifically, the front-end is typically responsible for interpreting sensor data and providing an initial estimate of the robot’s trajectory and the environment map; the back-end refines the initial estimate by performing global optimization to improve overall accuracy.
@@
Recent advancements in the field of SLAM have shown that supervised learning-based methods can deliver remarkable performance in front-end motion estimation (Wang et al., 2021b; Teed and Deng, 2021). These methods apply machine learning algorithms that depend on external supervision, usually in the form of a labeled dataset. Once trained, the model can make estimations without being explicitly programmed for the task. Meanwhile, geometry-based approaches remain crucial for the system’s back-end, primarily tasked with reducing front-end drift (Qin et al., 2018; Xu et al., 2025). These techniques employ geometrical optimization, such as pose graph optimization (PGO) (Labbe and Michaud, 2014), to maintain global consistency and minimize trajectory drift. However, their front-end and back-end components operate independently and are connected only in one direction. This means that the data-driven front-end model cannot receive feedback from the back-end for joint error correction. As a result, this decoupled approach might lead to suboptimal performance, adversely affecting the overall performance of the current SLAM systems.
@@
The choice of front-end odometry network and the back-end optimization are flexible in iSLAM. Here we illustrate one example of using a stereo visual-inertial odometry network as the front-end and a pose-velocity graph optimization as the back-end (Fu et al., 2024). Specifically, the front-end odometry consists of two independent efficient modules, i.e., a stereo odometry network and an IMU network (Qiu et al., 2024). The stereo odometry network is responsible for pose estimation, while the IMU network estimates both pose and velocity. These estimations are then incorporated as constraints (edges) in the back-end pose-velocity graph, linking the pose and velocity nodes. This setup allows for the optimization of pose and velocity nodes to reduce inconsistency between the visual and inertial measurements. As a result, the LL cost $L$ can be any estimation inconsistency between graph nodes:
@@ 4.5 Discrete Optimization
[…]
@@
where $\bm{z}_{i}$ is an instance among $S$ samplings. Note that $\nabla_{\theta}f(\bm{z}|\theta)$ is not a distribution function and thus (50c) is not an expectation and cannot be approximated by sampling. In contrast, (50e) is an expectation after the LDT is applied. The log-derivative estimator is a well-studied technique widely used in reinforcement learning (Sutton and Barto, 2018). The advantage is that it is unbiased but the drawback is that it often suffers from high variance (Grathwohl et al., 2017).
@@
where $\epsilon_{i}$ is an instance among $S$ samplings. Compared to the log-derivative estimator, the reparameterization estimator has a lower variance but sometimes introduces bias when model assumptions are not perfectly aligned with the true data, or if there are approximations in the model (Li et al., 2022a). Reparameterization trick has also been well studied and widely used in variational autoencoders (Doersch, 2016).
@@
We next introduce a simple but effective variance-reduced estimator based on control variate (Nelson, 1990), as defined in Lemma 2. It has been widely used in Monte Carlo estimation methods (Geffner and Domke, 2018; Richter et al., 2020; Leluc et al., 2022).
@@
Inspired by this control variate technique, we can construct a variance-reduced gradient estimator for $\nabla_{\bm{\theta}}g$ by designing a surrogate network (control variate) $s(\bm{z})$ approximating the reasoning process $g(\bm{z})$ (Grathwohl et al., 2017). Specifically, substitute ${\bm{\gamma}}=-1$ into (52), we have $X^{*}=X-Y+\beta$, this gives us the new control variate estimator $\hat{g}^{\prime}_{c}$:
@@
The multiple traveling salesman problem (MTSP) seeks tours for multiple agents such that all cities are visited exactly once while minimizing a cost function defined over the tours. MTSP is critical in many robotic applications where a team of robots is required to collectively visit a set of target locations, e.g., unmanned aerial vehicles (Sundar and Rathinam, 2013; Ma et al., 2019), automated agriculture (Carpio et al., 2021), warehouse logistics (Wurman et al., 2008; Ren et al., 2023b). The MTSP is renowned for its NP-hardness (Bektas, 2006). Intuitively, MTSP involves many decision variables including assigning the cities to the agents and determining the visiting order of the assigned cities. For example, an MTSP with $100$ cities and $10$ agents involves $\frac{100!\times 99!}{10!\times 89!}\approx 10^{20000}$ possible solutions, while the number of atoms in the observable universe is estimated to be “merely” $10^{78}$ to $10^{82}$ (Gaztañaga, 2023).
MTSP can be categorized by Min-Sum MTSP and Min-Max MTSP. Intuitively, Min-Sum MTSP aims to minimize the sum of tour lengths over all agents, while Min-Max MTSP aims to reduce the longest tour length. Though Min-Sum MTSP has been extensively studied (Bertazzi et al., 2015), Min-Max MTSP remains under-explored, and many techniques for Min-Sum MTSP may not directly apply to Min-Max MTSP. From the perspective of learning-based approaches, Min-Max MTSP is particularly challenging because the Min-Max operation is non-differentiable, which makes it harder to use gradient-based optimization and leads to a more complicated search for the global optimum. In this example, we will focus on the Min-Max MTSP, while our method may also apply to the Min-Sum MTSP. For the sake of brevity, we will use the abbreviation “MTSP” to refer to the Min-Max MTSP unless indicated otherwise.
@@
Due to the high complexity, classic MTSP solvers such as Google’s OR-Tools routing library meet difficulties for large-scale problems such as those with more than 500 cities (Furnon and Perron, 2023). To overcome this issue, there has been a notable shift towards employing neural network models to tackle MTSP (Hu et al., 2020; Xin et al., 2021; Miki et al., 2018; Vinyals et al., 2015; Kool et al., 2018; Nazari et al., 2018; Park et al., 2021; Ouyang et al., 2021; Khalil et al., 2017; Costa et al., 2020; Wu et al., 2021). However, these methods still have fundamental limitations, particularly in their ability to generalize to new, unseen situations, and in consistently finding high-quality solutions for large-scale problems. Supervised models struggle with limited supervision on small-scale problems and lack feasible supervision for large-scale instances, leading to poor generalization (Xin et al., 2021; Miki et al., 2018; Vinyals et al., 2015). Reinforcement learning (RL)-based approaches usually exploit implementations of the policy gradient algorithm, such as the Reinforce algorithm (Williams, 1992) and its variants. However, they often face the issue of slow convergence and highly sub-optimal solutions (Hu et al., 2020; Park et al., 2021). Other strategies like training greedy policy networks (Ouyang et al., 2021; Nazari et al., 2018; Kool et al., 2018; Khalil et al., 2017) and iteratively improving solutions (Costa et al., 2020; Wu et al., 2021) are also proposed but often get stuck at local optima.
@@
where $D$ computes the traveling cost given a tour, and $L$ returns the maximum route length among all agents. The UL optimization consists of two separate components: the UL cost $U_{\bm{\theta}}$ is to optimize the allocation network $f({\bm{\theta}})$, while $U_{\bm{\gamma}}$ is to optimize the surrogate network $s({\bm{\gamma}})$. Inspired by the control variant gradient estimator (56) and to have a lower gradient variance, we design the UL costs $U_{\bm{\theta}}$ and $U_{\bm{\gamma}}$ as

### v7

@@ 4.3 Constrained Optimization
We next illustrate the scenarios in which the LL cost $L$ in (1b) is subject to the general constraints in (1c). We discuss two cases with equality and inequality constraints, respectively. Constrained optimization is a thoroughly explored field, and related findings were presented in (dontchev2009implicit) and summarized in (gould2021deep). This study will focus on the integration of constrained optimization into our special form of BLO (1) under the framework of IL.
@@
Assume the LL cost $L(\cdot)$ and the constraint $\xi(\cdot)$ are 2nd-order differentiable near $({\bm{\theta}},{\bm{\gamma}},\bm{\phi}^{*}({\bm{\theta}},{\bm{\gamma}}))$ and the Hessian matrix $H$ below is invertible, we then have
@@
which means that the gradient of LL w.r.t. $\bm{\phi}$ has the same direction as the gradient of constraint $\xi$ w.r.t. $\bm{\phi}$ at the optimal solution $\bm{\mu}$$^{*}$ and $\bm{\nu}$$^{*}$. Taking the derivative of (40) w.r.t. $\bm{\psi}$ and noting that $\bm{\phi}^{*}$ and $\lambda$ are functions of $\bm{\psi}$, thus
@@
The analysis is similar to the equality case, with minor differences. Since the solution $\bm{\phi}^{*}$ with the inequality constraint (44) is a local minimum of the original inequality-constrained problem, it is reasonable to assume that the inequality constraints are active at $\bm{\phi}^{*}$. Assume $\bm{\phi}^{*}(\bm{\psi})$ exists, functions $L(\cdot)$ and $\xi(\cdot)$ are 2nd-order differentiable in the neighborhood of $(\bm{\psi},\bm{\phi}^{*}(\bm{\psi}))$, and the inequality constraint is active at $\bm{\phi}^{*}(\bm{\psi})$. Then, we can derive the same implicit derivative $\nabla\bm{\phi}^{*}(\bm{\psi})$ as in (38) of the equality case, except that the dual variable $\lambda\leq 0$ and $\nabla\bm{\phi}^{*}(\bm{\psi})$ needs to be computed using one-sided partial $\lambda=0$ due to non-differentiability, meaning we approach the point from only one direction (either left or right) to define the gradient, as the function may have abrupt changes or ‘kinks’ at that point.
@@
While we have derived closed-form solutions for the equality constraint in (35) and the inequality constraint in (44), calculating the matrix inversions $H^{-1}$ and $L_{\bm{\phi}}H^{-1}L_{\bm{\phi}}^{\top}]^{-1}$ can be computationally expensive. To mitigate this, one may employ the approach discussed in Section 4.1, which involves approximating the matrix-inverse-vector product $H^{-1}y$ by solving the linear system $\min_{v}\frac{1}{2}v^{\top}Hv-v^{\top}y$. This complicated computing process can sometimes be simplified by specific robot settings. We next present an example of robot control illustrating this process.
@@
When the Hessian matrices $\frac{\partial^{2}L}{\partial^{2}\phi}$ and $\frac{\partial^{2}\xi}{\partial^{2}\phi}$ of the LL function $L$ and the constraint function $\xi$ are invertible, we can guarantee the matrix $H$ in (38) is invertible. One sufficient condition for such invertibility is that LL and $\xi$ are strongly convex. Motivated by the approximate Karush-Kuhn-Tucker (KKT) condition (andreani2010new), which is characterized as an optimality condition for nonlinear program, we can use the residual function in (yao2024constrained, Eq. (15)) as an optimality measurement. As a result, the optimality can be established and guaranteed under these conditions.
@@
Modeling and control of nonlinear dynamics precisely is critical for robotic applications such as aerial manipulation (he2023image; geng2020cooperative) and off-road navigation (triest2023learning). Although classic methods based on physics have excellent generalization abilities, they rely heavily on problem-specific parameter tuning. Methods like state feedback (mellinger2011minimum) and optimal control (garcia1989model; ji2020robust) require precise system modeling to predict and control dynamics effectively. However, it is challenging to accurately model unpredictable factors such as wind gusts, boundary layer effects, uneven terrain, or hidden dynamics of chaotic nonlinear effects.
@@
In recent years, researchers have spent efforts to integrate physical modeling into data-driven methods and have gained remarkable success in the field of system control (o2022neural; hu2023off). For example, amos2018differentiable developed the differentiable model predictive control (MPC) to combine the physics-based modeling with data-driven methods, enabling learning dynamic models and control policies in an end-to-end manner (jin2020pontryagin). However, many prior studies depend on expert demonstrations or labeled data for supervised learning (jin2022learning; song2022policy). While supervised learning yields effective results in the training datasets, it suffers from challenging conditions such as unseen environments and external disturbances.
@@
This discrepancy between the predicted states and the IMU network measurement reflects the models’s imperfectness, indicating the learnable parameters $\bm{\theta}$} in both the network and the MPC module can be adjusted. Specifically, we adopt an IMU denoising model, AirIMU (qiu2023airimu) as the perception network, which is a neural model that can propagate system uncertainty through a time horizon. The differentiable MPC (Diff-MPC) from PyPose (wang2023pypose) is selected as the physics-based reasoning module.
@@ 4.4 Second-order Optimization
We next illustrate the scenario that LL cost $L$ in (1b) requires 2nd-order optimizers such as Gauss-Newton (GN) and Levenberg–Marquardt (LM). Because both GN and LM are iterative methods, one can leverage unrolled differentiation listed in Algorithm 1 to solve BLO. For the second strategy implicit differentiation, we could combine (31e) with a linear system solver in (32) in Section 4.2 to compute the implicit gradient $\nabla_{\bm{\theta}}U$ and $\nabla_{\bm{\gamma}}U$. Similar to the practical approach with LL first-order optimization, we can first use a second-order optimizer to solve the LL problem in (1b) to obtain approximates $\hat{\bm{\mu}}$ and $\hat{\bm{\nu}}$ of the solutions $\bm{\mu}$$^{*}$ and $\bm{\nu}$$^{*}$, which are then incorporated into (31e) to obtain an approximate $\widehat{\nabla}_{\bm{\theta}}U$ of the UL gradient $\nabla_{\bm{\theta}}U$ (similarly for $\widehat{\nabla}_{\bm{\gamma}}U$). Then, the UL gradient approximates $\widehat{\nabla}_{\bm{\theta}}U$ and $\widehat{\nabla}_{\bm{\gamma}}U$ are used to optimize the target variables $\bm{\theta}$} and $\bm{\gamma}$}. The first-order approximation without any second-order derivative computations as in (33) can be also developed for a practical training speedup.
@@
The theoretical guarantee is similar to that for the first-order methods in Section 4.2, as long as the second-order LL optimizers such as Gauss-Newton and LM approximate $\mu^{*}$ and $\nu^{*}$ to the target accuracy $\epsilon$. For example, if the LL problem is strongly convex, the second-order optimization can find an $\epsilon$-accurate solution with an accelerated linear convergence rate (ji2021lower, Theorem 9 and 10).
@@
Simultaneous localization and mapping (SLAM) is a critical technology in computer vision and robotics (aulinas2008slam; cadena2016past). It aims to simultaneously track the trajectory of a robot and build a map of the environment. SLAM is essential in diverse robotic applications, such as indoor navigation (wang2017non), underwater exploration (suresh2020active), and space missions (dor2021visual). A SLAM system generally adheres to a front-end and back-end architecture. Specifically, the front-end is typically responsible for interpreting sensor data and providing an initial estimate of the robot’s trajectory and the environment map; the back-end refines the initial estimate by performing global optimization to improve overall accuracy.
@@
Recent advancements in the field of SLAM have shown that supervised learning-based methods can deliver remarkable performance in front-end motion estimation (wang2021tartanvo; teed2021droid). These methods apply machine learning algorithms that depend on external supervision, usually in the form of a labeled dataset. Once trained, the model can make estimations without being explicitly programmed for the task. Meanwhile, geometry-based approaches remain crucial for the system’s back-end, primarily tasked with reducing front-end drift (qin2018vins; xu2025airslam). These techniques employ geometrical optimization, such as pose graph optimization (PGO) (pgo_labbe2014online), to maintain global consistency and minimize trajectory drift. However, their front-end and back-end components operate independently and are connected only in one direction. This means that the data-driven front-end model cannot receive feedback from the back-end for joint error correction. As a result, this decoupled approach might lead to suboptimal performance, adversely affecting the overall performance of the current SLAM systems.
@@
The choice of front-end odometry network and the back-end optimization are flexible in iSLAM. Here we illustrate one example of using a stereo visual-inertial odometry network as the front-end and a pose-velocity graph optimization as the back-end (fu2024islam). Specifically, the front-end odometry consists of two independent efficient modules, i.e., a stereo odometry network and an IMU network (qiu2023airimu). The stereo odometry network is responsible for pose estimation, while the IMU network estimates both pose and velocity. These estimations are then incorporated as constraints (edges) in the back-end pose-velocity graph, linking the pose and velocity nodes. This setup allows for the optimization of pose and velocity nodes to reduce inconsistency between the visual and inertial measurements. As a result, the LL cost $L$ can be any estimation inconsistency between graph nodes:
@@ 4.5 Discrete Optimization
$=\nabla_{\theta}\int f(\bm{z}|\theta)g(\bm{z})d\bm{z}$ (51b)
$=\int\nabla_{\theta}f(\bm{z}|\theta)g(\bm{z})d\bm{z}\quad\text{(Leibniz rule)}$ (51c)
$=\int f(\bm{z}|\theta)\frac{\nabla_{\theta}f(\bm{z}|\theta)}{f(\bm{z}|\theta)}g(\bm{z})d\bm{z}\quad\left(\text{Multiply }\frac{f(\bm{z}|\theta)}{f(\bm{z}|\theta)}\right)$ (51d)
$=\int f(\bm{z}|\theta)\nabla_{\theta}\log f(\bm{z}|\theta)g(\bm{z})d\bm{z}\quad\text{(The LDT)}$ (51e)
@@
where $\bm{z}_{i}$ is an instance among $S$ samplings. Note that $\nabla_{\theta}f(\bm{z}|\theta)$ is not a distribution function and thus (51c) is not an expectation and cannot be approximated by sampling. In contrast, (51e) is an expectation after the LDT is applied. The log-derivative estimator is a well-studied technique widely used in reinforcement learning (sutton2018reinforcement). The advantage is that it is unbiased but the drawback is that it often suffers from high variance (grathwohl2017backpropagation).
@@
where $\epsilon_{i}$ is an instance among $S$ samplings. Compared to the log-derivative estimator, the reparameterization estimator has a lower variance but sometimes introduces bias when model assumptions are not perfectly aligned with the true data, or if there are approximations in the model (li2022implicit). Reparameterization trick has also been well studied and widely used in variational autoencoders (doersch2016tutorial).
@@
We next introduce a simple but effective variance-reduced estimator based on control variate (nelson1990control), as defined in Lemma 2. It has been widely used in Monte Carlo estimation methods (geffner2018using; richter2020vargrad; leluc2022quadrature).
@@
Inspired by this control variate technique, we can construct a variance-reduced gradient estimator for $\nabla_{\bm{\theta}}g$ by designing a surrogate network (control variate) $s(\bm{z})$ approximating the reasoning process $g(\bm{z})$ (grathwohl2017backpropagation). Specifically, substitute ${\bm{\gamma}}=-1$ into (53), we have $X^{*}=X-Y+\beta$, this gives us the new control variate estimator $\hat{g}^{\prime}_{c}$:
@@
The multiple traveling salesman problem (MTSP) seeks tours for multiple agents such that all cities are visited exactly once while minimizing a cost function defined over the tours. MTSP is critical in many robotic applications where a team of robots is required to collectively visit a set of target locations, e.g., unmanned aerial vehicles (sundar2013algorithms; ma2019coordinated), automated agriculture (carpio2021mp), warehouse logistics (wurman2008coordinating; 10109784). The MTSP is renowned for its NP-hardness (bektas2006multiple). Intuitively, MTSP involves many decision variables including assigning the cities to the agents and determining the visiting order of the assigned cities. For example, an MTSP with $100$ cities and $10$ agents involves $\frac{100!\times 99!}{10!\times 89!}\approx 10^{20000}$ possible solutions, while the number of atoms in the observable universe is estimated to be “merely” $10^{78}$ to $10^{82}$ (gaztanaga2023mass).
MTSP can be categorized by Min-Sum MTSP and Min-Max MTSP. Intuitively, Min-Sum MTSP aims to minimize the sum of tour lengths over all agents, while Min-Max MTSP aims to reduce the longest tour length. Though Min-Sum MTSP has been extensively studied (bertazzi2015min), Min-Max MTSP remains under-explored, and many techniques for Min-Sum MTSP may not directly apply to Min-Max MTSP. From the perspective of learning-based approaches, Min-Max MTSP is particularly challenging because the Min-Max operation is non-differentiable, which makes it harder to use gradient-based optimization and leads to a more complicated search for the global optimum. In this example, we will focus on the Min-Max MTSP, while our method may also apply to the Min-Sum MTSP. For the sake of brevity, we will use the abbreviation “MTSP” to refer to the Min-Max MTSP unless indicated otherwise.
@@
Due to the high complexity, classic MTSP solvers such as Google’s OR-Tools routing library meet difficulties for large-scale problems such as those with more than 500 cities (ortools_routing). To overcome this issue, there has been a notable shift towards employing neural network models to tackle MTSP (hu2020reinforcement; xin2021neurolkh; miki2018applying; vinyals2015pointer; kool2018attention; nazari2018reinforcement; park2021schedulenet; ouyang2021improving; khalil2017learning; d2020learning; wu2021learning). However, these methods still have fundamental limitations, particularly in their ability to generalize to new, unseen situations, and in consistently finding high-quality solutions for large-scale problems. Supervised models struggle with limited supervision on small-scale problems and lack feasible supervision for large-scale instances, leading to poor generalization (xin2021neurolkh; miki2018applying; vinyals2015pointer). Reinforcement learning (RL)-based approaches usually exploit implementations of the policy gradient algorithm, such as the Reinforce algorithm (williams1992simple) and its variants. However, they often face the issue of slow convergence and highly sub-optimal solutions (hu2020reinforcement; park2021schedulenet). Other strategies like training greedy policy networks (ouyang2021improving; nazari2018reinforcement; kool2018attention; khalil2017learning) and iteratively improving solutions (d2020learning; wu2021learning) are also proposed but often get stuck at local optima.
@@
where $D$ computes the traveling cost given a tour, and LL returns the maximum route length among all agents. The UL optimization consists of two separate components: the UL cost $U_{\bm{\theta}}$ is to optimize the allocation network $f({\bm{\theta}})$, while $U_{\bm{\gamma}}$ is to optimize the surrogate network $s({\bm{\gamma}})$. Inspired by the control variant gradient estimator (57) and to have a lower gradient variance, we design the UL costs $U_{\bm{\theta}}$ and $U_{\bm{\gamma}}$ as